OpenStudy (anonymous):

What improper fraction is equal to 2 1/2

terenzreignz (terenzreignz):

Given a mixed-number, it is equal to this corresponding improper fraction... \[\huge a\frac{p}q=\frac{q\times a+p}{q}\]

OpenStudy (anonymous):

To turn a mixed number into an improper fraction : Take the whole number and multiply it by the denominator, then add that to the numerator, then put that over the same denominator it already had. For example ... 4 2/3 (whole number 4 times 3 = 12 plus 2 = 14 over original denominator 3 = 14/3) You understand ?